# Structure-Based Molecule Optimization via Gradient-Guided Bayesian Update

### v2

@@ 4.1 Equivariant Guidance for multi-modality Molecular Data
In the continuous domain, guidance over $\bm{\theta}$ (i.e. $\bm{\mu}$) is analogous to guided diffusions, since guiding $\bm{\mu}$ corresponds to guiding noisy latent $\mathbf{y}$ using the gradient over it:
$\bm{\mu}_{i}^{*}=\bm{\mu}_{i}+\sigma\mathbf{g}_{\bm{\mu}}=\frac{\bm{\mu}_{i-1}\rho_{i-1}+(\mathbf{y}^{x}+\sigma\frac{\rho_{i}}{\alpha_{i}}\mathbf{g}_{\bm{\mu}})\alpha_{i}}{\rho_{i}}=\frac{\bm{\mu}_{i-1}\rho_{i-1}+(\mathbf{y}^{x}+\sigma\mathbf{g}_{\mathbf{y}^{x}})\alpha_{i}}{\rho_{i}}$ (8)
where by chain rule $\mathbf{g}_{\mathbf{y}^{x}}=-\nabla_{\mathbf{y}^{x}}E(\bm{\theta},t_{i})|_{\bm{\theta}=\bm{\theta}_{i-1}}=\mathbf{g}_{\bm{\mu}}\frac{\partial h}{\partial\mathbf{y}^{x}}=\mathbf{g}_{\bm{\mu}}\frac{\alpha_{i}}{\rho_{i}}$. This builds a connection between the guidance over $\bm{\theta}$ and over latent $\mathbf{y}$ for continuous data.
Guiding $\mathbf{z}$ for Discrete $\mathbf{v}$.
For $N$-dimensional discrete types $\mathbf{v}\in\{1,\dots,K\}^{N}$, the noisy latent represents the counts of each type among $K$ types, where we have $\mathbf{y}^{v}\sim\mathcal{N}(\mathbf{y}^{v}|\alpha^{\prime}(K\mathbf{e}_{\mathbf{v}}-\mathbb{1}),\alpha^{\prime}K\mathbb{I})$, $\mathbf{e}_{\mathbf{v}}=[\mathbf{e}_{\mathbf{v}^{(1)}},\dots,\mathbf{e}_{\mathbf{v}^{(N)}}]\in\mathbb{R}^{KN}$, $\mathbf{e}_{\mathbf{v}^{(j)}}=\delta_{\mathbf{v}^{(j)}}\in\mathbb{R}^{K}$ with Kronecker delta function $\delta$. This is mathematically grounded with further explanation in Appendix B.
The aggregated $\mathbf{z}_{i}$ as a posterior belief is updated given the prior $\mathbf{z}_{0}=\frac{\mathbb{1}}{\mathbb{K}}$ as:
$h(\mathbf{z}_{i-1},\mathbf{y}^{v},\alpha^{\prime}_{i})$ $=\frac{\exp(\mathbf{y}^{v})\mathbf{z}_{i-1}}{\sum_{k=1}^{K}\exp(\mathbf{y}^{v}_{k})(\mathbf{z}_{i-1})_{k}}$ (9)
where the redundant $\alpha^{\prime}_{i}=\beta^{\prime}_{i}-\beta^{\prime}_{i-1}$ with $\beta^{\prime}_{i}=\beta^{\prime}_{1}(\frac{i}{n})^{2}$, given a positive hyperparameter $\beta^{\prime}_{1}$.
@@
In the discrete domain, guiding all latents $\mathbf{y}$ amounts to a reweight of the Categorical distribution for $\bm{\theta}$ (i.e. $\mathbf{z}$), changing the probability of each class in accordance with the gradient. Take an extreme case to illustrate, where $\mathbf{g}_{\mathbf{y}^{v}}$ is filled with one-hot vectors $\delta_{d}$:
$(\mathbf{z}^{*}_{i})_{k}=\frac{\exp(\mathbf{y}^{v}_{k})(\mathbf{z}_{i-1})_{k}}{\sum_{k^{\prime}=1}^{K}\exp(\mathbf{y}^{v}_{k^{\prime}})(\mathbf{z}_{i-1})_{k^{\prime}}+[\exp(\sigma^{\prime})-1]\exp(\mathbf{y}^{v}_{d})(\mathbf{z}_{i-1})_{d}}=(\mathbf{z}_{i})_{k}\frac{1}{1+C}<(\mathbf{z}_{i})_{k}$ (10)
for all $k\neq d$, where $C=[\exp(\sigma^{\prime})-1]\exp(\mathbf{y}^{v}_{d})(\mathbf{z}_{i-1})_{d}/[{\sum_{k^{\prime}=1}^{K}\exp(\mathbf{y}^{v}_{k^{\prime}})}]>0$ as the variance $\sigma^{\prime}>0$. It is obvious that the guidance lowers the probability for all classes but the favored $d$, redistributing the mass for discrete data in a more structured way than diffusion counterparts.
@@
Our proposed guided sampling that utilizes joint gradient signals is still equivariant as shown in the proposition below, with proof in Appendix C.2.
@@
4.2 Guided Bayesian Update With Backward Correction
Algorithm 1 Gradient Guided Sampling of MolJO with Backward Correction
0:  network $\bm{\Phi}(\bm{\theta},t,\mathbf{p})$, noise schedules $[\beta(t),\beta^{\prime}(t)]$, number of sample steps $n$, back correction steps $k$, number of atom types $K$, energy function $E(\bm{\theta},\mathbf{p},t)$, guidance scale $s$
1:  Initialize belief $\bm{\theta}:=[\bm{\mu},\mathbf{z}]\leftarrow[\mathbf{0},\mathbf{\frac{1}{K}}]$triangleright Continuous aggregations for different modalities
@@ 4.2 Guided Bayesian Update With Backward Correction
3:     $[t,t_{-k}]\leftarrow[\frac{i-1}{n},\max(0,\frac{i-k-1}{n})]$triangleright Discrete timesteps in $[0,1)$
4:     $[\hat{\mathbf{x}},\hat{\mathbf{e}}_{\mathbf{v}}]\leftarrow\bm{\Phi}(\bm{\theta},t,\mathbf{p})$triangleright Network predicts clean data
5:     $[\mathbf{g}_{\bm{\mu}},\mathbf{g}_{\mathbf{y}}]\leftarrow[-\nabla_{\bm{\mu}}E(\bm{\theta},\mathbf{p},t),-\nabla_{\mathbf{y}}E(\bm{\theta},\mathbf{p},t)]$triangleright Gradient guidance
6:     $[\rho,\rho_{-k},\Delta\beta,\Delta\beta^{\prime}]\leftarrow[1+\beta(t),1+\beta(t_{-k}),\beta(t)-\beta(t_{-k}),\beta^{\prime}(t)-\beta^{\prime}(t_{-k})]$
7:     Retrieve $\bm{\mu}_{-k},\mathbf{z}_{-k}$ from the past $\triangleright$ Backward correction starting point
8:     Sample $\bm{\mu}\sim\mathcal{N}([\frac{\rho(\Delta\beta\hat{\mathbf{x}}+\bm{\mu}_{-k}\rho_{-k})+s\Delta\beta\mathbf{g}_{\bm{\mu}}}{\rho^{2}},\frac{\Delta\beta}{\rho^{2}}\mathbf{I})$triangleright Guidance for continuous $\mathbf{x}$
9:     Sample $\mathbf{y}\sim\mathcal{N}(\Delta\beta^{\prime}(t)(K\hat{\mathbf{e}}_{\mathbf{v}}-\mathbf{1}+sK\mathbf{g}_{\mathbf{y}}),\Delta\beta^{\prime}(t)K\mathbf{I})$triangleright Guidance for discrete $\mathbf{v}$
10:     Update $\mathbf{z}\leftarrow\frac{\exp(\mathbf{y})\mathbf{z}_{-k}}{\sum_{i=1}^{K}\exp(\mathbf{y}_{i})(\mathbf{z}_{-k})_{i}}$triangleright Aggregate latents for discrete data
11:  end for
12:  $[\hat{\mathbf{x}},\hat{\mathbf{e}}_{\mathbf{v}}]\leftarrow\bm{\Phi}(\bm{\theta},1,\mathbf{p})$triangleright Network predicts final output
13:  $\hat{\mathbf{v}}\leftarrow\arg\max(\hat{\mathbf{e}}_{\mathbf{v}})$triangleright Final discrete output from p.m.f.
14:  return  $[\hat{\mathbf{x}},\hat{\mathbf{v}}]$
Here we propose a general backward correction sampling strategy inspired from the optimization perspective, and analyze its effect on aligning the gradients. Recall that from Eq. 1 we can aggregate $\bm{\theta}_{i}$ based on latents from the previous step:
@@
[…]
[3pt]↓
[…]
@@
and we arrive at the $k-1$ step corrected estimation of $p_{\phi}$:
$p_{\phi}(\bm{\theta}_{n}|\bm{\theta}_{n-1},\bm{\theta}_{n-k})$ $=\underset{p_{O}(\hat{\mathbf{x}}_{n}|\bm{\Phi}(\bm{\theta}_{n-1},t_{n}))}{\mathbb{E}}p_{U}(\bm{\theta}_{n}|\bm{\theta}_{n-k},\hat{\mathbf{x}}_{n};\sum_{i=n-k+1}^{n}\alpha_{i})$ (13)
Plugging Eq. 7 and 9 together with the sender distributions defined above into the right hand side according to Eq. 1, yields the analytic form of the backward corrected Bayesian update
$p_{U}(\bm{\mu}_{n}|\bm{\mu}_{n-k},\hat{\mathbf{x}}_{n})$ $=\mathcal{N}\Big{(}\frac{\Delta\beta\hat{\mathbf{x}}_{n}+\bm{\mu}_{n-k}\rho_{n-k}}{\rho_{n}},\frac{\Delta\beta}{\rho_{n}^{2}}\mathbb{I}\Big{)}$ (14)
$p_{U}(\mathbf{z}_{n}|\mathbf{z}_{n-k},\hat{\mathbf{v}}_{n})$ $=\underset{\mathbf{y}\sim\mathcal{N}(\mathbf{y}|\Delta\beta^{\prime}(K\mathbf{e}_{\hat{\mathbf{v}}_{n}}-\mathbb{1}),\Delta\beta^{\prime}K\mathbb{I})}{\mathbb{E}}\delta\Big{(}\mathbf{z}_{n}-\frac{\exp(\mathbf{y})\mathbf{z}_{n-k}}{\sum_{i=1}^{K}\exp(\mathbf{y}_{i})(\mathbf{z}_{n-k})_{i}}\Big{)}$ (15)
where $\hat{\mathbf{m}}=[\hat{\mathbf{x}},\hat{\mathbf{v}}]$ is drawn from the output distribution $p_{O}({\hat{\mathbf{m}}}~{}|~{}\bm{\Phi}(\bm{\theta}_{i-1},t_{i-1},\mathbf{p}))$ given pocket $\mathbf{p}$, $\Delta\beta=\beta_{n}-\beta_{n-k}$ and $\Delta\beta^{\prime}=\beta^{\prime}_{n}-\beta^{\prime}_{n-k}$ are obtained from corresponding schedules.
Figure 2: Gradient similarity, where $k$ denotes backward correction window size. For $2001<k<$ 200, the similarity before timestep $k$ is omitted for it overlaps with $k=200$, i.e. covering all the past.
The concept of sliding window unifies different sampling strategies proposed by Graves et al. (2023) (k=1) and Qu et al. (2024) (k=n). To understand its effect, we visualize the cosine similarity of gradients at each step w.r.t. the previous step in Figure 2. By changing the size $k$ of sliding window, it succeeds in balancing sample quality (explore) and optimization efficiency (exploit), where it first focuses on exploring the molecular space with rapidly changing structures and gradients, and then exploits better-aligned guidance signals over gradually refined structures, achieving the best success rate as later shown.
In practice, we employ the gradient scale $s$ as a temperature parameter, which is equivalent to adopting $p_{E}^{s}(\bm{\theta},t)\propto\exp{[-sE(\bm{\theta},t)]}$. The general sampling procedure is summarized in Algorithm 1.
@@ 5.1 Experimental Setup
Following previous SBDD works (Luo et al., 2021; Peng et al., 2022; Guan et al., 2022), we utilize CrossDocked2020 (Francoeur et al., 2020) to train and test our model, and adopt the same processing that filters out poses with RMSD $>$ 1Å and clusters proteins based on 30% sequence identity, yielding 100,000 training poses and 100 test proteins.
@@
We divide all baselines into the following: (1) Generative models (Gen.), including AR (Luo et al., 2021), GraphBP (Liu et al., 2022), Pocket2Mol (Peng et al., 2022), FLAG (Zhang et al., 2023), DiffSBDD (Schneuing et al., 2022), TargetDiff (Guan et al., 2022), DecompDiff (Guan et al., 2023), IPDiff (Huang et al., 2024) and MolCRAFT (Qu et al., 2024), (2) Oracle-based optimization (Oracle Opt.) that rely on docking simulation in each round, such as RGA (Fu et al., 2022), and DecompOpt (Zhou et al., 2024), (3) Gradient-guided (Grad. Opt.) TAGMol (Dorna et al., 2024). Detailed descriptions of baselines are left in Appendix F.
@@
We employ the commonly used metrics as follows: (1) Affinity metrics calculated by Autodock Vina (Eberhardt et al., 2021), in which Vina Score calculates the raw energy of the given molecular pose residing in the pocket, Vina Min conducts a quick local energy minimization and scores the minimized pose, and Vina Dock performs a relatively longer search for optimal pose to calculate the lowest energy. Success Rate measures the percentage of generated molecules that pass certain criteria (Vina Dock $<$ -8.18, QED $>$ 0.25, SA $>$ 0.59) following Long et al. (2022) and Guan et al. (2022). (2) Molecular properties, including drug-likeness (QED) and synthesizability score (SA). (3) Metrics for sample distribution, such as diversity (Div). A more comprehensive set of metrics are detailed in Appendix F.

### v3

@@ 4.1 Equivariant Guidance for multi-modality Molecular Data
In the continuous domain, guidance over $\bm{\theta}^{x}$ is analogous to guided diffusions, since guiding $\bm{\theta}^{x}$ corresponds to guiding noisy latent $\mathbf{y}^{x}$ using the uncertainty-adjusted gradient $(\frac{\rho_{i}}{\alpha_{i}})^{2}\mathbf{g}_{\mathbf{y}^{x}}$ that accounts for how changes in the sample space $\mathbf{y}^{x}$ propagate to parameter space $\bm{\theta}^{x}$:
$\bm{\theta}^{x}_{i}$ $=\tilde{\bm{\theta}}^{x}_{i}+\sigma^{x}\mathbf{g}_{\bm{\theta}^{x}}$
$=\frac{\bm{\theta}^{x}_{i-1}\rho_{i-1}+(\tilde{\mathbf{y}}^{x}+\sigma^{x}\frac{\rho_{i}}{\alpha_{i}}\mathbf{g}_{\bm{\theta}^{x}})\alpha_{i}}{\rho_{i}}$
$=\frac{\bm{\theta}^{x}_{i-1}\rho_{i-1}+(\tilde{\mathbf{y}}^{x}+\sigma^{x}(\frac{\rho_{i}}{\alpha_{i}})^{2}\mathbf{g}_{\mathbf{y}^{x}})\alpha_{i}}{\rho_{i}}$ (7)
Guiding $\bm{\theta}^{v}$ for Discrete $\mathbf{v}$.
For $N$-dimensional discrete types $\mathbf{v}\in\{1,\dots,K\}^{N}$, the noisy latent represents the counts of each type among $K$ types, where we have $\mathbf{y}^{v}\sim\mathcal{N}(\mathbf{y}^{v}|\alpha(K\mathbf{e}_{\mathbf{v}}-\mathbb{1}),\alpha K\mathbb{I})$, $\mathbf{e}_{\mathbf{v}}=[\mathbf{e}_{\mathbf{v}^{(1)}},\dots,\mathbf{e}_{\mathbf{v}^{(N)}}]\in\mathbb{R}^{KN}$, $\mathbf{e}_{\mathbf{v}^{(j)}}=\delta_{\mathbf{v}^{(j)}}\in\mathbb{R}^{K}$ with Kronecker delta function $\delta$. Further explanation is left to Appendix B.
$\bm{\theta}^{v}_{i}$ as a posterior belief is updated from the prior $\bm{\theta}^{v}_{0}=\frac{\mathbb{1}}{\mathbb{K}}$:
$h(\bm{\theta}^{v}_{i-1},\mathbf{y}^{v},\alpha_{i})$ $=\frac{\exp(\mathbf{y}^{v})\bm{\theta}^{v}_{i-1}}{\sum_{k=1}^{K}\exp(\mathbf{y}^{v}_{k})(\bm{\theta}^{v}_{i-1})_{k}}$ (8)
where the redundant $\alpha_{i}=\beta^{v}_{i}-\beta^{v}_{i-1}$ with $\beta^{v}_{i}=\beta^{v}_{1}(\frac{i}{n})^{2}$, given a positive hyperparameter $\beta^{v}_{1}$.
@@
In the discrete domain, guiding all latents $\mathbf{y}^{v}$ amounts to a reweight of the Categorical distribution for $\bm{\theta}^{v}$, changing the probability of each class in accordance with the gradient. Take an extreme case to illustrate, where $\mathbf{g}_{\mathbf{y}^{v}}$ is filled with one-hot vectors $\delta_{d}$:
$(\bm{\theta}^{v}_{i})_{k}$ $=\frac{\exp(\tilde{\mathbf{y}}^{v}_{k})(\bm{\theta}^{v}_{i-1})_{k}}{\sum_{l}\exp(\tilde{\mathbf{y}}^{v}_{l})(\bm{\theta}^{v}_{i-1})_{l}+[\exp(\sigma^{v})-1]\exp(\tilde{\mathbf{y}}^{v}_{d})(\bm{\theta}^{v}_{i-1})_{d}}$
$=(\tilde{\bm{\theta}}^{v}_{i})_{k}\frac{1}{1+C}<(\tilde{\bm{\theta}}^{v}_{i})_{k}$ (9)
for all $k\neq d$, where $C=\frac{[\exp(\sigma^{v})-1]\exp(\tilde{\mathbf{y}}^{v}_{d})({\bm{\theta}}^{v}_{i-1})_{d}}{[{\sum_{l=1}^{K}\exp(\tilde{\mathbf{y}}^{v}_{l})}]}>0$ as the variance $\sigma^{v}>0$. It is obvious that the guidance lowers the probability for all classes but the favored $d$, redistributing the mass for discrete data in a more structured way than diffusion counterparts.
@@
Our proposed guided sampling that utilizes joint gradient signals is still equivariant as shown in the proposition below, with the proof in Appendix C.2.
@@
The guided sampling process preserves SE(3)-equivariance when $\bm{\Phi}$ is SE(3)-equivariant, if the energy function $E(\bm{\theta},\mathbf{p},t)$ is also SE(3)-equivariant and the complex is shifted to the space where the protein’s Center of Mass (CoM) is zero.
4.2 Bayesian Update With Backward Correction
Algorithm 1 Gradient Guided Sampling of MolJO
0:  network $\bm{\Phi}(\bm{\theta},t,\mathbf{p})$, schedules $[\beta^{x}(t),\beta^{v}(t)]$, number of sample steps $n$, back correction steps $k$, number of atom types $K$, energy function $E(\bm{\theta},\mathbf{p},t)$, guidance scale $s$
1:  Initialize belief $\bm{\theta}:=[\bm{\theta}^{x},\bm{\theta}^{v}]\leftarrow[\mathbf{0},\mathbf{\frac{1}{K}}]$
@@ 4.2 Bayesian Update With Backward Correction
3:     $[t,t_{-k}]\leftarrow[\frac{i-1}{n},\max(0,\frac{i-k-1}{n})]$
4:     $[\hat{\mathbf{x}},\hat{\mathbf{e}}_{\mathbf{v}}]\leftarrow\bm{\Phi}(\bm{\theta},t,\mathbf{p})$
5:     $[\mathbf{g}_{\bm{\theta}^{x}},\mathbf{g}_{\bm{\theta}^{v}}]\leftarrow[-\nabla_{\bm{\theta}^{x}}E(\bm{\theta},\mathbf{p},t),-\nabla_{\bm{\theta}^{v}}E(\bm{\theta},\mathbf{p},t)]$
6:     $[\rho,\rho_{-k},\Delta\beta^{x},\Delta\beta^{v}]\leftarrow[1+\beta^{x}(t),1+\beta^{x}(t_{-k}),\beta^{x}(t)-\beta^{x}(t_{-k}),\beta^{v}(t)-\beta^{v}(t_{-k})]$
7:     Retrieve $\bm{\theta}^{x}_{-k},\bm{\theta}^{v}_{-k}$ from the past
8:     Sample $\bm{\theta}^{x}$ according to Eq. 4 and 13
9:     Sample $\mathbf{y}^{v}$ and update $\bm{\theta}^{v}$ according to Eq. 5 and 4.2
10:  end for
11:  $[\hat{\mathbf{x}},\hat{\mathbf{e}}_{\mathbf{v}}]\leftarrow\bm{\Phi}(\bm{\theta},1,\mathbf{p})$
12:  $\hat{\mathbf{v}}\leftarrow\arg\max(\hat{\mathbf{e}}_{\mathbf{v}})$
13:  return $[\hat{\mathbf{x}},\hat{\mathbf{v}}]$
Here we propose a general backward correction sampling strategy inspired from the optimization perspective, and analyze its effect on aligning the gradients. Recall that from Eq. 1 we can aggregate $\bm{\theta}_{i}$ from previous latents:
@@
$p_{\phi}(\bm{\theta}_{i}~{}|~{}\bm{\theta}_{i-1},\bm{\theta}_{i-2})=$
$\underset{p_{O}(\hat{\mathbf{x}}_{i}|\bm{\Phi}(\bm{\theta}_{i-1},t_{i}))}{\mathbb{E}}{\underset{p_{U}(\bm{\theta}_{i-1}|\bm{\theta}_{i-2},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\hat{\mathbf{x}}_{i}}};\alpha_{i-1})}{\mathbb{E}}}$ $p_{U}(\bm{\theta}_{i}~{}|~{}\bm{\theta}_{i-1},\hat{\mathbf{x}}_{i};\alpha_{i})$
@@
where the original $\hat{\mathbf{x}}_{i-1}\sim p_{O}(\hat{\mathbf{x}}_{i-1}|\bm{\Phi}(\bm{\theta}_{i-2},t_{i-1}))$ that has been used to update $\bm{\theta}_{i-1}$ from $\bm{\theta}_{i-2}$ at previous step, is now replaced by the optimized $\hat{\mathbf{x}}_{i}$. By iteratively tracing back, we arrive at the $k-1$ step corrected estimation of $p_{\phi}$:
$p_{\phi}(\bm{\theta}_{n}|\bm{\theta}_{n-1},\bm{\theta}_{n-k})$ $=$
$\underset{p_{O}(\hat{\mathbf{x}}_{n}|\bm{\Phi}(\bm{\theta}_{n-1},t_{n}))}{\mathbb{E}}$ $p_{U}(\bm{\theta}_{n}|\bm{\theta}_{n-k},\hat{\mathbf{x}}_{n};\sum_{i=n-k+1}^{n}\alpha_{i})$ (12)
Plugging Eq. 6 and 8 together with the sender distributions defined above into the right hand side according to Eq. 1, yields the form of the backward corrected Bayesian update
$p_{U}(\bm{\theta}^{x}_{n}|\bm{\theta}^{x}_{n-k},\hat{\mathbf{x}}_{n})$ $=\mathcal{N}\Big{(}\frac{\Delta\beta\hat{\mathbf{x}}_{n}+\bm{\theta}^{x}_{n-k}\rho_{n-k}}{\rho_{n}},\frac{\Delta\beta}{\rho_{n}^{2}}\mathbb{I}\Big{)}$ (13)
$p_{U}(\bm{\theta}^{v}_{n}|\bm{\theta}^{v}_{n-k},\hat{\mathbf{v}}_{n})$ $=\underset{\mathbf{y}\sim\mathcal{N}(\mathbf{y}|\Delta\beta^{v}(K\mathbf{e}_{\hat{\mathbf{v}}_{n}}-\mathbb{1}),\Delta\beta^{v}K\mathbb{I})}{\mathbb{E}}$
$\delta\Big{(}\bm{\theta}^{v}_{n}-\frac{\exp(\mathbf{y})\bm{\theta}^{v}_{n-k}}{\sum_{i=1}^{K}\exp(\mathbf{y}_{i})(\bm{\theta}^{v}_{n-k})_{i}}\Big{)}$ (14)
where $\hat{\mathbf{m}}=[\hat{\mathbf{x}},\hat{\mathbf{v}}]$ is drawn from the output distribution $p_{O}({\hat{\mathbf{m}}}|\bm{\Phi}(\bm{\theta}_{i-1},t_{i-1},\mathbf{p}))$ given pocket $\mathbf{p}$, $\Delta\beta^{x}=\beta^{x}_{n}-\beta^{x}_{n-k}$ and $\Delta\beta^{v}=\beta^{v}_{n}-\beta^{v}_{n-k}$ are obtained from corresponding accuracy schedules. The concept of sliding window
Figure 2: Gradient cosine similarity, where $k$ denotes the backward correction window size. For $2001<k<$ 200, the similarity before timestep $k$ is omitted for it overlaps with $k=200$, i.e. covering all the past.
unifies different sampling strategies proposed by Graves et al. (2023) (k=1) and Qu et al. (2024) (k=n). To understand its effect, we visualize the cosine similarity of gradients at each step w.r.t. the previous step in Fig. 2. By changing the size $k$ of sliding window, it succeeds in balancing sample quality (explore) and optimization efficiency (exploit), where it first focuses on exploring the molecular space with rapidly changing structures and gradients, and then exploits better-aligned guidance signals over gradually refined structures.
In practice, we employ the gradient scale $s$ as a temperature parameter, equivalent to $p_{E}^{s}(\bm{\theta},t)\propto\exp{[-sE(\bm{\theta},t)]}$. We further bypass the derivative $\frac{\partial\bm{\theta}^{v}}{\partial\mathbf{y}^{v}}=\bm{\theta}^{v}(1-\bm{\theta}^{v})$ to stabilize the gradient flow. The general sampling procedure is summarized in Algorithm 1.
@@ 5.1 Experimental Setup
Following previous SBDD works (Luo et al., 2021), we utilize CrossDocked2020 (Francoeur et al., 2020) to train and test our model, and adopt the same processing that filters out poses with RMSD $>$ 1Å and clusters proteins based on 30% sequence identity, yielding 100,000 training poses and 100 test proteins.
@@
We divide all baselines into the following: (1) Generative models (Gen), including AR (Luo et al., 2021), GraphBP (Liu et al., 2022), Pocket2Mol (Peng et al., 2022), FLAG (Zhang et al., 2023), DiffSBDD (Schneuing et al., 2022), TargetDiff (Guan et al., 2022), DecompDiff (Guan et al., 2023), IPDiff (Huang et al., 2024) and MolCRAFT (Qu et al., 2024), (2) Oracle-based optimization (Oracle) that rely on docking simulation in each round, such as AutoGrow4 (Spiegel & Durrant, 2020), RGA (Fu et al., 2022), and DecompOpt (Zhou et al., 2024a), (3) Gradient-guided (Grad) TAGMol (Dorna et al., 2024). Detailed descriptions of baselines are left in Appendix F.
@@
We employ the commonly used metrics as follows: (1) Affinity metrics calculated by Autodock Vina (Eberhardt et al., 2021), in which Vina Score calculates the raw energy of the given molecular pose residing in the pocket, Vina Min conducts a quick local energy minimization and scores the minimized pose, and Vina Dock performs a relatively longer search for optimal pose to calculate the lowest energy. Success Rate measures the percentage of generated molecules that pass certain criteria (Vina Dock $<$ -8.18, QED $>$ 0.25, SA $>$ 0.59) following Guan et al. (2022). (2) Molecular properties, including drug-likeness (QED) and synthesizability score (SA). (3) Metrics for sample distribution, such as diversity (Div). A more comprehensive set of metrics are detailed in Appendix F.